OpenStudy (janu16):

what is the approximate surface area of the sphere? C=37.68

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ c}}ircumference=2\pi r\qquad \{{\color{brown}{ c}}=37.68\qquad {\color{brown}{ 37.68}}=2\pi r\) solve for "r", see what you get for the radius and use that to find the area of the circle

OpenStudy (janu16):

How do you find the radius

OpenStudy (triciaal):

does it matter that this is not a circle it is a sphere? surface area of a sphere = 4 * pi* r^2

OpenStudy (janu16):

I actually found the answer but thank you!

OpenStudy (triciaal):